OpenStudy (anonymous):

Find the approximate value of b in the triangle shown. 13 12 8 5

OpenStudy (anonymous):

use cosine law, have an idea what is it ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

\(b^2=a^2+c^2-2ac~ \cos B\)

OpenStudy (anonymous):

b2=12^2+10^2−2ac cosB ?

OpenStudy (anonymous):

maybe u know it by its name "law of cosines"

OpenStudy (anonymous):

no i really dont:( dont really have had time to learn it

OpenStudy (anonymous):

its ok :) just memorize the formula and everything will be ok ur on the right track , u need to sub a and c as well like this http://prntscr.com/5uqic6

OpenStudy (anonymous):

here u can see this link also http://www.mathsisfun.com/algebra/trig-cosine-law.html

OpenStudy (anonymous):

i got 4 from it...

OpenStudy (anonymous):

:O how?

OpenStudy (anonymous):

12^2+10^2 - 2*12*10 cos 24 = 142

OpenStudy (anonymous):

how do i get it be one of my options?

OpenStudy (anonymous):

wait \(b^2= 12^2+10^2 - 2*12*10 cos 24 =24.74.....\) \(b=\sqrt(24.74.....)=4.97...\) almost 5 when u said 4 like it was the right answer but i waned to be sure about approximation value

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

eh np